OpenStudy (cindy):

its complicated let me rite it in words and see if you understand better one plus a fraction and the fraction is one / two plus one third

OpenStudy (angela210793):

is it...

OpenStudy (angela210793):

1+1/(2+1/3)?

OpenStudy (cindy):

ca you showme th steps

OpenStudy (angela210793):

yea ofc but is it this.....1 + 1 ------ 2+ 1/3

OpenStudy (cindy):

yes it is

OpenStudy (angela210793):

first let's transmute 1 1 1 3 ------=-----------=---------=----- 2+1/3 (2*3+1)/3 7/3 7 then 1+3/7=7/7+3/7=10/7

OpenStudy (angela210793):

Do u understand it now?

OpenStudy (cindy):

nope..sorry

OpenStudy (angela210793):

Cindy please look carefully .....

OpenStudy (angela210793):

1st i transmuted 1 ------ 2+ 1/3 and the result was 3/7 then i summed it with 1 as the question asks it

OpenStudy (cindy):

but theres another 1+ infront of the raction

OpenStudy (angela210793):

yes i know.....1+3/7=7/7+3/7=10/7 ...that is the 1 i have summed with 3/7

OpenStudy (angela210793):

take a look at this...maybe u'll understand :)

OpenStudy (cindy):

1 + 1 2 + 1/3 ITS LIKE TIS BUT THRES A LINE BETWEEN THE 1 AND 2+ 1/3

OpenStudy (angela210793):

i know Cindy...this is wht i have solved....

OpenStudy (angela210793):

firstly i solved 1 -------- 2 + 1/3 and then i added 1